OpenStudy (anonymous):

Francisco has a savings account balance of $2,033.88. The interest rate on the account is 2.9% compounded monthly. If he opened the account nine years ago, what was the value of his initial deposit?

OpenStudy (anonymous):

first draw the cash flow diagram

OpenStudy (kropot72):

The formula to solve this is \[A=P(1+\frac{r}{n})^{tn}\] where A is the amount after t years, P is the initial deposit, r is the annual interest rate as a decimal and n is the number of compounding events per year. Plugging in values we get \[2033.88=P(1+\frac{0.029}{12})^{9\times 12}=P \times 1.0024167^{108}\] Therefore wse can find P from \[P=\frac{2033.88}{1.0024167^{108}}=you\ can\ calculate\]